OpenStudy (anonymous):

A laptop computer is purchased for $1300. After each year, the resale value decreases by 25% . What will the resale value be after 4 years?

OpenStudy (dtan5457):

You can go year by year and subtract Year 1: 1400-(.25x1400)=1050 Year 2:1050-(.25x1050)=787.5 Year 3: 787.5-(.25x787.5)=590.625 now just do one more year

OpenStudy (anonymous):

Thank you so much!

OpenStudy (dtan5457):

you got the last year?

OpenStudy (anonymous):

443?

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

yeah, that's about right

OpenStudy (anonymous):

it says no :(